# Diffusion Bridge Implicit Models

### v3

@@ Appendix A Related Work
Fast sampling of diffusion models can be classified into training-free and training-based methods. A prevalent training-free fast sampler is the denoising diffusion implicit models (DDIMs) (Song et al., 2021a) that employ alternative non-Markovian generation processes in place of DDPMs, a discrete-time diffusion model. ScoreSDE (Song et al., 2021c) further links discrete-time DDPMs to continuous-time score-based models, unrevealing the generation process to be ordinary and stochastic differential equations (ODEs and SDEs). Subsequent training-free samplers concentrate on developing dedicated numerical solvers to the diffusion ODE or SDE, particularly Heun’s methods (Karras et al., 2022) and exponential integrators (Zhang & Chen, 2022; Lu et al., 2022b; Zheng et al., 2023a; Gonzalez et al., 2023). These methods typically require around 10 steps for high-quality generation. In contrast, training-based methods, particularly adversarial distillation (Sauer et al., 2023) and consistency distillation (Song et al., 2023b; Kim et al., 2023), become notable for their ability to achieve high-quality generation with just one or two steps. Developing dedicated differential equation solvers for DDBMs, or exploring bridge distillation methods, would be promising avenues for future research.
@@
Diffusion bridges (De Bortoli et al., 2021; Liu et al., 2023; Somnath et al., 2023; Zhou et al., 2023; Chen et al., 2023; Shi et al., 2024) are a promising generative variant of diffusion models for modeling the transport between two arbitrary distributions. One line of work is the diffusion Schrodinger bridge models (De Bortoli et al., 2021; Shi et al., 2024), which solves an entropy-regularized optimal transport problem between two probability distributions. However, their reliance on expensive iterative procedures has limited their application scope, particularly for high-dimensional data. Subsequent works have endeavored to enhance the tractability of the Schrodinger bridge problem by making assumptions such as paired data (Liu et al., 2023; Somnath et al., 2023; Chen et al., 2023). On the other hand, DDBMs (Zhou et al., 2023) construct diffusion bridges via Doob’s $h$-transform, offering a reverse-time perspective of a diffusion process conditioned on given endpoints. This approach aligns the design spaces and training algorithms of DDBMs closely with those of score-based generative models, leading to state-of-the-art performance in image translation tasks. However, the sampling procedure of DDBMs still relies on inefficient simulations of differential equations, lacking theoretical insights to develop efficient samplers.
@@ Appendix B Proofs
$\mathrm{d}\bm{x}_{t}=\left[f(t)\bm{x}_{t}-g^{2}(t)\left(\frac{1}{2}\nabla_{\bm{x}_{t}}\log q(\bm{x}_{t}|\bm{x}_{T})-\nabla_{\bm{x}_{t}}\log q_{tT}(\bm{x}_{T}|\bm{x}_{t})\right)\right]\mathrm{d}t$ (33)
with the data predictor $\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$. We replace the bridge score $\nabla_{\bm{x}_{t}}\log q(\bm{x}_{t}|\bm{x}_{T})$ with the network $\bm{s}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$, which is related to $\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$ by Eqn. (14). Besides, $\nabla_{\bm{x}_{t}}\log q_{tT}(\bm{x}_{T}|\bm{x}_{t})$ can be analytically computed as
$\nabla_{\bm{x}_{t}}\log q_{tT}(\bm{x}_{T}|\bm{x}_{t})$ $=\nabla_{\bm{x}_{t}}\log\frac{q(\bm{x}_{t}|\bm{x}_{0},\bm{x}_{T})q(\bm{x}_{T}|\bm{x}_{0})}{q(\bm{x}_{t}|\bm{x}_{0})}$ (34)
@@ C.1 Markov Property of the Generalized Diffusion Bridges
Due to the marginal preservation property (Proposition 3.1), we have $q^{(\rho)}(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{T})=q(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{T})$ and $q^{(\rho)}(\bm{x}_{t_{n}}|\bm{x}_{0},\bm{x}_{T})=q(\bm{x}_{t_{n}}|\bm{x}_{0},\bm{x}_{T})$, where $q(\bm{x}_{t}|\bm{x}_{0},\bm{x}_{T})=\mathcal{N}(a_{t}\bm{x}_{T}+b_{t}\bm{x}_{0},c_{t}^{2}\bm{I})$ is the forward transition kernel in Eqn. (6). To identify whether $q^{(\rho)}(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{t_{n}},\bm{x}_{T})$ is Markovian, we only need to examine the dependence of $\bm{x}_{t_{n+1}}$ on $\bm{x}_{0}$. To this end, we proceed to derive conditions under which $\nabla_{\bm{x}_{t_{n+1}}}\log q^{(\rho)}(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{T})$ involves terms concerning $\bm{x}_{0}$.
Specifically, $\nabla_{\bm{x}_{t_{n+1}}}\log q^{(\rho)}(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{T})$ can be calculated as
$\nabla_{\bm{x}_{t_{n+1}}}\log q^{(\rho)}(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{T})$ (50)
@@
$=$ $-\frac{\sqrt{c_{t_{n}}^{2}-\rho_{n}^{2}}(a_{t_{n}}\bm{x}_{T}+b_{t_{n}}\bm{x}_{0}+\sqrt{c^{2}_{t_{n}}-\rho^{2}_{n}}\frac{\bm{x}_{t_{n+1}}-a_{t_{n+1}}\bm{x}_{T}-b_{t_{n+1}}\bm{x}_{0}}{c_{t_{n+1}}})}{c_{t_{n+1}}\rho_{n}^{2}}$
@@ C.3 Special Cases and Relationship with Prior Works
When the noise schedule $\alpha_{t}=1,T=1$ and $\sigma_{t}=\sqrt{\beta t}$, the forward process becomes $q(\bm{x}_{t}|\bm{x}_{0},\bm{x}_{1})=\mathcal{N}(t\bm{x}_{T}+(1-t)\bm{x}_{0},\beta t(1-t))$ which is a Brownian bridge. When $\beta\rightarrow 0$, there will be no intermediate noise and the forward process is similar to flow matching (Lipman et al., 2022). In this limit, the DBIM ($\eta=1$) updating rule from time $t$ to time $s<t$ will become $\bm{x}_{s}=s\bm{x}_{T}+(1-s)\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})=\frac{s}{t}\bm{x}_{t}+(1-\frac{s}{t})\bm{x}_{\theta}(\bm{x}_{t},t)=\bm{x}_{t}-(t-s)\bm{v}_{\theta}(\bm{x}_{t},t)$. Here we define $\bm{v}_{\theta}(\bm{x}_{t},t)\coloneqq\frac{\bm{x}_{t}-\bm{x}_{\theta}(\bm{x}_{t},t)}{t}$ as the velocity function of the probability flow (i.e., the drift of the ODE) in flow matching methods. Therefore, in the flow matching case, DBIM is a simple Euler step of the flow.
@@
The previous work I2SB (Liu et al., 2023) also employs diffusion bridges with discrete timesteps, though their noise schedule is restricted to the variance exploding (VE) type with $f(t)=0$ in the forward process. For generation, they adopt a similar approach to DDPM (Ho et al., 2020) by iterative sampling from the posterior distribution $p_{\theta}(\bm{x}_{n-1}|\bm{x}_{n})$, which is a parameterized and shortened diffusion bridge between the endpoints $\hat{\bm{x}}_{0}=\bm{x}_{\theta}(\bm{x}_{n},t_{n},\bm{x}_{N})$ and $\bm{x}_{n}$. Since the posterior distribution is not conditioned on $\bm{x}_{T}$ (except through the parameterized network), the corresponding forward diffusion bridge is Markovian. Thus, the posterior sampling in I2SB is a special case of DBIM by setting $\eta=0$ and $f(t)=0$.
@@ F.1 Model Details
For the image inpainting task on ImageNet 256$\times$256 with 128$\times$128 center mask, DDBMs do not provide available checkpoints. Therefore, we train a new model from scratch using the noise schedule of I2SB (Liu et al., 2023). The network is initialized from the pretrained class-conditional diffusion model on ImageNet 256$\times$256 provided by (Dhariwal & Nichol, 2021), while additionally conditioned on $\bm{x}_{T}$. The data prediction model in this case is parameterized by the network $\bm{F}_{\theta}$ as $\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})=\bm{x}_{t}-\sigma_{t}\bm{F}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$ and trained by minimizing the loss $\mathcal{L}(\theta)=\mathbb{E}_{t,\bm{x}_{0},\bm{x}_{T}}\left[\frac{1}{\sigma_{t}^{2}}\|\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})-\bm{x}_{0}\|_{2}^{2}\right]$. We train the model on 8 NVIDIA A800 GPU cards with a batch size of 256 for 400k iterations, which takes around 19 days.
@@ F.3 License
I2SB https://github.com/NVlabs/I2SB Liu et al. (2023) CC-BY-NC-SA-4.0
@@
Appendix G Additional Samples

### v4

@@ Appendix A Related Work
Fast sampling of diffusion models can be classified into training-free and training-based methods. A prevalent training-free fast sampler is the denoising diffusion implicit models (DDIMs) (Song et al., 2021a) that employ alternative non-Markovian generation processes in place of DDPMs, a discrete-time diffusion model. ScoreSDE (Song et al., 2021c) further links discrete-time DDPMs to continuous-time score-based models, unrevealing the generation process to be ordinary and stochastic differential equations (ODEs and SDEs). DDIM can be generalized to develop integrators for broader diffusion models (Zhang et al., 2022; Pandey et al., 2023). The concept of implicit sampling, in a broad sense, can also be extended to discrete diffusion models (Chen et al., 2024; Zheng et al., 2024), although there are fundamental differences in their underlying mechanisms. Subsequent training-free samplers concentrate on developing dedicated numerical solvers to the diffusion ODE or SDE, particularly Heun’s methods (Karras et al., 2022) and exponential integrators (Zhang & Chen, 2022; Lu et al., 2022b; Zheng et al., 2023a; Gonzalez et al., 2023). These methods typically require around 10 steps for high-quality generation. In contrast, training-based methods, particularly adversarial distillation (Sauer et al., 2023) and consistency distillation (Song et al., 2023b; Kim et al., 2023), become notable for their ability to achieve high-quality generation with just one or two steps. Our work serves as a thorough exploration of training-free fast sampling of DDBMs. Exploring bridge distillation methods, such as consistency bridge distillation (He et al., 2024), would be promising future research avenues to decrease the inference cost further.
@@
Diffusion bridges (De Bortoli et al., 2021; Chen et al., 2021b; Liu et al., 2023b; Somnath et al., 2023; Zhou et al., 2023; Chen et al., 2023; Shi et al., 2024; Deng et al., 2024) are a promising generative variant of diffusion models for modeling the transport between two arbitrary distributions. One line of work is the diffusion Schrodinger bridge models (De Bortoli et al., 2021; Chen et al., 2021b; Shi et al., 2024; Deng et al., 2024), which solves an entropy-regularized optimal transport problem between two probability distributions. However, their reliance on expensive iterative procedures has limited their application scope, particularly for high-dimensional data. Subsequent works have endeavored to enhance the tractability of the Schrodinger bridge problem by making assumptions such as paired data (Liu et al., 2023b; Somnath et al., 2023; Chen et al., 2023). On the other hand, DDBMs (Zhou et al., 2023) construct diffusion bridges via Doob’s $h$-transform, offering a reverse-time perspective of a diffusion process conditioned on given endpoints. This approach aligns the design spaces and training algorithms of DDBMs closely with those of score-based generative models, leading to state-of-the-art performance in image translation tasks. However, the sampling procedure of DDBMs still relies on inefficient simulations of differential equations, lacking theoretical insights to develop efficient samplers. BBDM (Li et al., 2023) and I3SB (Wang et al., 2024a) extend the concept of DDIM to the contexts of Brownian bridge and I2SB (Liu et al., 2023b), respectively. SinSR (Wang et al., 2024b) is also motivated by DDIM, while the application is concentrated on the mean-reverting diffusion process, which ends in a Gaussian instead of a delta distribution. In contrast to them, our work provides the first systematic exploration of implicit sampling within the broader DDBM framework, offering theoretical insights and connections while also proposing novel high-order diffusion bridge solvers.
@@ Appendix B Proofs
$\mathrm{d}\bm{x}_{t}=\left[f(t)\bm{x}_{t}-g^{2}(t)\left(\frac{1}{2}\nabla_{\bm{x}_{t}}\log q(\bm{x}_{t}|\bm{x}_{T})-\nabla_{\bm{x}_{t}}\log q_{T|t}(\bm{x}_{T}|\bm{x}_{t})\right)\right]\mathrm{d}t$ (33)
with the data predictor $\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$. We replace the bridge score $\nabla_{\bm{x}_{t}}\log q(\bm{x}_{t}|\bm{x}_{T})$ with the network $\bm{s}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$, which is related to $\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$ by Eqn. (14). Besides, $\nabla_{\bm{x}_{t}}\log q_{T|t}(\bm{x}_{T}|\bm{x}_{t})$ can be analytically computed as
$\nabla_{\bm{x}_{t}}\log q_{T|t}(\bm{x}_{T}|\bm{x}_{t})$ $=\nabla_{\bm{x}_{t}}\log\frac{q(\bm{x}_{t}|\bm{x}_{0},\bm{x}_{T})q(\bm{x}_{T}|\bm{x}_{0})}{q(\bm{x}_{t}|\bm{x}_{0})}$ (34)
@@ C.1 Markov Property of the Generalized Diffusion Bridges
Due to the marginal preservation property (Proposition 3.1), we have $q^{(\rho)}(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{T})=q(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{T})$ and $q^{(\rho)}(\bm{x}_{t_{n}}|\bm{x}_{0},\bm{x}_{T})=q(\bm{x}_{t_{n}}|\bm{x}_{0},\bm{x}_{T})$, where $q(\bm{x}_{t}|\bm{x}_{0},\bm{x}_{T})=\mathcal{N}(a_{t}\bm{x}_{T}+b_{t}\bm{x}_{0},c_{t}^{2}\bm{I})$ is the forward transition kernel in Eqn. (6). To identify whether $q^{(\rho)}(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{t_{n}},\bm{x}_{T})$ is Markovian, we only need to examine the dependence of $\bm{x}_{t_{n+1}}$ on $\bm{x}_{0}$. To this end, we proceed to derive conditions under which $\nabla_{\bm{x}_{t_{n+1}}}\log q^{(\rho)}(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{t_{n}},\bm{x}_{T})$ involves terms concerning $\bm{x}_{0}$.
Specifically, $\nabla_{\bm{x}_{t_{n+1}}}\log q^{(\rho)}(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{t_{n}},\bm{x}_{T})$ can be calculated as
$\nabla_{\bm{x}_{t_{n+1}}}\log q^{(\rho)}(\bm{x}_{t_{n+1}}|\bm{x}_{0},\bm{x}_{t_{n}},\bm{x}_{T})$ (50)
@@
$=$ $-\frac{\sqrt{c_{t_{n}}^{2}-\rho_{n}^{2}}(a_{t_{n}}\bm{x}_{T}+b_{t_{n}}\bm{x}_{0}+\sqrt{c^{2}_{t_{n}}-\rho^{2}_{n}}\frac{\bm{x}_{t_{n+1}}-a_{t_{n+1}}\bm{x}_{T}-b_{t_{n+1}}\bm{x}_{0}}{c_{t_{n+1}}}-\bm{x}_{t_{n}})}{c_{t_{n+1}}\rho_{n}^{2}}$
@@ C.3 Special Cases and Relationship with Prior Works
When the noise schedule $\alpha_{t}=1,T=1$ and $\sigma_{t}=\sqrt{\beta t}$, the forward process becomes $q(\bm{x}_{t}|\bm{x}_{0},\bm{x}_{1})=\mathcal{N}(t\bm{x}_{T}+(1-t)\bm{x}_{0},\beta t(1-t))$ which is a Brownian bridge. When $\beta\rightarrow 0$, there will be no intermediate noise and the forward process is similar to flow matching (Lipman et al., 2022; Albergo et al., 2023). In this limit, the DBIM ($\eta=1$) updating rule from time $t$ to time $s<t$ will become $\bm{x}_{s}=s\bm{x}_{T}+(1-s)\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})=\frac{s}{t}\bm{x}_{t}+(1-\frac{s}{t})\bm{x}_{\theta}(\bm{x}_{t},t)=\bm{x}_{t}-(t-s)\bm{v}_{\theta}(\bm{x}_{t},t)$. Here we define $\bm{v}_{\theta}(\bm{x}_{t},t)\coloneqq\frac{\bm{x}_{t}-\bm{x}_{\theta}(\bm{x}_{t},t)}{t}$ as the velocity function of the probability flow (i.e., the drift of the ODE) in flow matching methods. Therefore, in the flow matching case, DBIM is a simple Euler step of the flow.
@@
The previous work I2SB (Liu et al., 2023b) also employs diffusion bridges with discrete timesteps, though their noise schedule is restricted to the variance exploding (VE) type with $f(t)=0$ in the forward process. For generation, they adopt a similar approach to DDPM (Ho et al., 2020) by iterative sampling from the posterior distribution $p_{\theta}(\bm{x}_{n-1}|\bm{x}_{n})$, which is a parameterized and shortened diffusion bridge between the endpoints $\hat{\bm{x}}_{0}=\bm{x}_{\theta}(\bm{x}_{n},t_{n},\bm{x}_{N})$ and $\bm{x}_{n}$. Since the posterior distribution is not conditioned on $\bm{x}_{T}$ (except through the parameterized network), the corresponding forward diffusion bridge is Markovian. Thus, the posterior sampling in I2SB is a special case of DBIM by setting $\eta=0$ and $f(t)=0$.
@@ F.1 Model Details
For the image inpainting task on ImageNet 256$\times$256 with 128$\times$128 center mask, DDBMs do not provide available checkpoints. Therefore, we train a new model from scratch using the noise schedule of I2SB (Liu et al., 2023b). The network is initialized from the pretrained class-conditional diffusion model on ImageNet 256$\times$256 provided by (Dhariwal & Nichol, 2021), while additionally conditioned on $\bm{x}_{T}$. The data prediction model in this case is parameterized by the network $\bm{F}_{\theta}$ as $\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})=\bm{x}_{t}-\sigma_{t}\bm{F}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$ and trained by minimizing the loss $\mathcal{L}(\theta)=\mathbb{E}_{t,\bm{x}_{0},\bm{x}_{T}}\left[\frac{1}{\sigma_{t}^{2}}\|\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})-\bm{x}_{0}\|_{2}^{2}\right]$. We train the model on 8 NVIDIA A800 GPU cards with a batch size of 256 for 400k iterations, which takes around 19 days.
@@ F.3 License
I2SB https://github.com/NVlabs/I2SB Liu et al. (2023b) CC-BY-NC-SA-4.0
@@
Appendix G Additional Results
G.1 Runtime Comparison
Table 8 shows the inference time of DBIM and previous methods on a single NVIDIA A100 under different settings. We use torch.cuda.Event and torch.cuda.synchronize to accurately compute the runtime. We evaluate the runtime on 8 batches (dropping the first batch since it contains extra initializations) and report the mean and std. We can see that the runtime is proportional to NFE. This is because the main computation costs are the serial evaluations of the large neural network, and the calculation of other coefficients requires neglectable costs. Therefore, the speedup for the NFE is approximately the actual speedup of the inference time.
Table 8: Runtime of different methods to generate a single batch (second / batch, $\pm$±std) on a single NVIDIA A100, varying the number of function evaluations (NFE).
Method NFE
5 10 15 20
Center $128\times 128$ Inpainting, ImageNet $256\times 256$ (batch size = 16)
I2SB (Liu et al., 2023b) $2.8128\pm 0.0111$ $5.6049\pm 0.0152$ $8.3919\pm 0.0166$ $11.1494\pm 0.0259$
DDBM (Zhou et al., 2023) $2.8711\pm 0.0318$ $5.7283\pm 0.0572$ $8.3787\pm 0.1667$ $11.0678\pm 0.3061$
DBIM ($\eta=0$) $2.8755\pm 0.0706$ $5.7810\pm 0.1494$ $8.5890\pm 0.2730$ $11.1613\pm 0.3372$
DBIM (2nd-order) $2.8859\pm 0.0675$ $5.7884\pm 0.1734$ $8.6284\pm 0.1907$ $11.5898\pm 0.2260$
DBIM (3rd-order) $2.9234\pm 0.0361$ $5.8109\pm 0.2982$ $8.6449\pm 0.2118$ $11.3710\pm 0.3237$
G.2 Diversity Score
We measure the diversity score (Batzolis et al., 2021; Li et al., 2023) on the ImageNet center inpainting task. We calculate the standard deviation of 5 generated samples (numerical range $0\sim 255$) given each observation (condition) $\bm{x}_{T}$, averaged over all pixels and 1000 conditions.
[…]
Method NFE
5 10 20 50 100 200 500
I2SB (Liu et al., 2023b) 3.27 4.45 5.21 5.75 5.95 6.04 6.15
DDBM (Zhou et al., 2023) - - 2.96 4.03 4.69 5.29 5.83
DBIM $\eta=0$ 3.74 4.56 5.20 5.80 6.10 6.29 6.42
$\eta=1$ 2.62 3.40 4.18 5.01 5.45 5.81 6.16
As shown in Table 9, the diversity score keeps increasing with larger NFE. DBIM ($\eta=0$) consistently surpasses the flow matching baseline I2SB, and DDBM’s hybrid sampler which introduces diversity through SDE steps. Surprisingly, we find that the DBIM $\eta=0$ case exhibits a larger diversity score than the $\eta=1$ case. This demonstrates that the booting noise can introduce enough stochasticity to ensure diverse generation. Moreover, the $\eta=0$ case tends to generate sharper images, which may favor the diversity score measured by pixel-level variance.
Appendix H Additional Samples